The half-life of an isotope is 2.0 hours. How much of a 50.0 g sample is left after 6.0 hours?


1
Expert's answer
2021-12-07T05:27:03-0500

The half-life of a radioactive isotope is 2.0 hours, i.e, "t_{1\/2}" = 2.0 hours.

Therefore, the radiactive decay constant is "λ = \frac{0.693}{t_{1/2}} = \frac{0.693}{2.0 \;hours}"

Let N0 = 50.0 g the initial amount of the radioisotope and Nt is the amount remaining after t = 6.0 hours.

Use the radioactive decay law.

"N(t) = N_0 \times e^{-λt} \\

ln (\frac{N_t }{N_0}) = -λ \times t"

Plug in values and get

"ln (\frac{N_t }{50.0}) = -\frac{0.693}{2.0 \;hours} \times 6.0 \; hours = -2.079 \\

\frac{N_t }{50.0} = e^{-2.079} \\

\frac{N_t }{50.0} = 0.12505 \\

N_t = 6.2525 \; g"

The amount of the radioactive isotope left is 6.2525 g (ans).